ASIAX:

Kim Barney pays a $290.00 annual premium for an insurance plan with a $500 deductible. The company pays 80% of the remaining expense. If Kim had $2,500.00 in medical expenses, calculate the following: a. Kim's annual premium plus deductible $____ b. The insurance company's payment of the medical expense. $____ c. Kim's total medical expenses divided by her premium plus deductible (to nearest hundredth). ____ d. Kim's medical expense less the insurance payment. $____ e. The percentage of her medical expense covered by insurance. % ____

Angle:

I think you can figure (a) out right?

Angle:

then for (b) the insurance only pays 80% of "the remaining expense" the "remaining expense" = 2500 - (deductible) = 2500 - 500 = 2000 insurance pays = 2000*0.8 = 1600

ASIAX:

right (a) is 790

ASIAX:

@Ultrilliam can you help me finish this?

Ultrilliam:

I'm unfortunately a bit preoccupied ATM; perhaps @Angle @Vocaloid could help you when they have time v.v

ASIAX:

ok thanks

ASIAX:

@Angle can you help me with part D & E?

Angle:

D wants to know how much does Kim have to pay herself after all the insurance and stuff which would be 2500 - 500 = 2000 - 1600 = 400 E ... 2500 - 400 = insurance covered that number divided by 2500 = percent of how much the insurance covered